Circular No. 7288

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999el IN NGC 6951 L. Cao, Y. L. Qiu, Q. Y. Qiao, and J. Y. Hu, Beijing Astronomical

Observatory (BAO), on behalf of the BAO Supernova Survey, report the discovery of an apparent supernova (mag ∼ 15.4) in NGC 6951 on an unfiltered CCD frame taken on Oct. 20.45 UT with a 0.60-m telescope at Xinglong station; an image taken on Oct. 21.44 yields mag ∼ 15.0 (there is a knot at the location of SN 1999el on the Digital Sky Survey image and on previous BAO images, and the magnitude estimates include the brightness of the supernova and the knot). SN 1999el is located at α = 20^{h}37^{m}17^{s}.83, δ = +66^{o}06′11′′.5 (equinox 2000.0), which is 21′′.8 east and 8′′.4 south of the nucleus of NGC 6951. Nothing appears on images taken on Oct. 19.2 by KAIT to limiting mag 18.0 (according to W. Li, University of California at Berkeley) or on Oct. 19.44 at BAO. Li adds that on Oct. 22.2, SN 1999el appeared at mag 15.2 on a KAIT image, and they find the supernova to be 21′′.7 east and 8′′.6 south of the galaxy nucleus. A KAIT image taken on 1998 Sept. 13 (limiting mag ∼ 20.5), shows two nearby foreground stars: one (mag ∼ 17.3) located ∼ 23′′.9 east and 8′′.4 south of the nucleus, and another (mag ∼ 18.6) located ∼ 19′′.4 east and 8′′.8 south of the nucleus. A. Filippenko (Berkeley) has identified a BAO spectrum of SN 1999el as that of a very early type-IIn supernova.

COMET C/1995 O1 (HALE-BOPP) A. Pearce, Nedlands, W. Australia, reports that visual observations

with a 0.41-m reflector on Oct. 20.84 and 21.84 UT show a bright 14thmag condensation that was not visible on Oct. 18, though m_{1} has not changed much. CCD observations by I. P. Griffin and M. Bos, obtained with the Auckland Observatory 0.5-m telescope on Oct. 15 and 22, also show a significantly more condensed coma on the second night.

Ephemeris extension to IAUC 7252:

1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 29 7^{h}04.^{m}2 −74^{o}32.^{′}8  9.765  9.670 81.^{o}6 5.^{o}8 12.8
Nov. 8 6 56.5 −75 34.1  9.851  9.743 80.9 5.8 12.9
18 6 45.5 −76 27.5  9.937  9.816 80.2 5.7 12.9
28 6 31.1 −77 10.5 10.020  9.889 79.5 5.6 13.0
Dec. 8 6 13.9 −77 40.7 10.102  9.961 79.0 5.6 13.0
18 5 54.9 −77 56.3 10.180 10.034 78.7 5.5 13.1
28 5 35.6 −77 56.9 10.256 10.106 78.5 5.5 13.1
1999 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green